OpenStudy (anonymous):

Can't figure out how to solve (factoring quadratic equations): 4x^2 +7x = 2

OpenStudy (anonymous):

first make it equal to 0 by making the equationb 4x^2+7x-2 =0. then use teh quadratic formula \[x = -b + or - \sqrt{b ^{2}-4ac}/ 2a\]

OpenStudy (anonymous):

Factored Wrong... The answer would end up being (x-2) (4x-1)=0 or x=2 and x=1/4

OpenStudy (anonymous):

Would you be able to outline step by step - not really understanding this at all. Issue seems to be when the 'c' is less than 'b' - can't figure out how to solve. THANK YOU!

OpenStudy (anonymous):

Set equal to zero: 4x^2=7x-2=0 Plug the standard form equation into the quadratic formula: \[-7\pm \sqrt{7^2-4(4)(-2)}/2(8)\] Simplify: \[-7\pm \sqrt{49+32}/8\] \[-7\pm \sqrt{81}/8\] \[(-7\pm9)/8\] Solve for X: \[(-7+9)/8=.25 or 1/4\] and \[(-7-9)/8= -2\]